OpenStudy (anonymous):

The length of a rectangle exceeds its width by 3 inches, and the area is 54 square inches. What are the length and width of the rectangle?

OpenStudy (anonymous):

l=w+3

OpenStudy (anonymous):

length=width+3 make sense right?

OpenStudy (anonymous):

length* width= area w *( w+3 ) =54

OpenStudy (anonymous):

w^2+3w=54

OpenStudy (anonymous):

w^2+3w-54 (w+9)(w-6)=0 w is either -9 or 6, but width can't be negative number so it is 6 length is 3 more than width so it is 9